OpenStudy (anonymous):

Find three solutions of the equation. 644-14-07-00-00_files/i0270000.jpg a. (–2, 3), (1, –3), (2, –4) c. (1, –3), (0, –1), (–1, 0) b. (–2, 3), (1, –3), (0, –1) d. (0, –1), (3, –9), (–2, 3)

OpenStudy (anonymous):

Where is the equation?

OpenStudy (anonymous):

y = -2x -1 @desiswag

OpenStudy (precal):

just test the points

OpenStudy (anonymous):

That's true you could do that @amandalepore

OpenStudy (precal):

(-2,3) this means when x=-2 y = 3 y=-2x-1 3=-2(-2)-1 are both sides equal? If so then that is a solution, if not then go to choice B and test those points

OpenStudy (anonymous):

so its B?

OpenStudy (precal):

no I said go test those points to see if it is B

OpenStudy (anonymous):

Whats the answer? :/

OpenStudy (anonymous):

Try to test the points

OpenStudy (anonymous):

I don't know how lol

OpenStudy (precal):

Welcome to open study. I know this is your first question. Our goal at this site it to help you understand the process (oh that means study)

OpenStudy (precal):

I just gave you an example

OpenStudy (precal):

(-2,3) is a coordinate (x,y) is the coordinate this is how we say x=-2 and y=3

OpenStudy (precal):

(-2,3) this means when x=-2 y = 3 y=-2x-1 3=-2(-2)-1 use the order of operation to evaluate the right hand side

OpenStudy (anonymous):

is it C?

OpenStudy (precal):

sorry we are not allowed to take quizzes, or exams for people

OpenStudy (anonymous):

No, I did the work I just want to know If im correct?

OpenStudy (precal):

perhaps @desiswag can explain it a little better than I can

OpenStudy (anonymous):

@amandalepore did both sides equal each other when you plugged in the points? if so, then c is the answer. when you plugged in the points, did both sides of the equation equal each other?

OpenStudy (anonymous):

Yes! so its C?

OpenStudy (anonymous):

c (the second choice listed) doesnt look quite right. Look at the last set of points, (-1,0). 0 = -2(-1) -1 --> 0 = 2 - 1 --> 0 =1

OpenStudy (anonymous):

try the third choice listed

OpenStudy (anonymous):

Right..I did some work , does C seem right?

OpenStudy (anonymous):

@desiswag

OpenStudy (anonymous):

c as in the third choice listed?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Yup! That's right! My bad, I thought c was the second choice in your question

OpenStudy (anonymous):

Thank you so much! life saver!

OpenStudy (anonymous):

yw :)